OpenStudy (anonymous):

Given the sequence 6, 18, 54, 162, ..., which expression would give the fifteenth term?.

OpenStudy (anonymous):

first term is 6, second term is 3x6, third it 3^2x6, so 14th is 3^13x6

OpenStudy (anonymous):

\[6*3^{14}=28697814 \]

OpenStudy (anonymous):

oh sorry 15th term is just 3^(14)x6

OpenStudy (anonymous):

\[T_n=6 \times 3^{n-1} \]

OpenStudy (anonymous):

foolfor the nth term gets a medal

OpenStudy (anonymous):

Tn=ar^n-1

OpenStudy (anonymous):

n=position of the term r=common ratio a=value of the first term

OpenStudy (anonymous):

It is a geometric progression

OpenStudy (anonymous):

Answer is 28697814